# Efficient Model-Stealing Attacks Against Inductive Graph Neural Networks

### v2

@@ 8 Conclusions
In this work, we introduce a new unsupervised GNN model-stealing attack. Only a single work [30] targeting inductive GNNs existed previously. Our proposed approach combines a contrastive learning objective with graph-specific augmentations, improving the efficient utilization of information derived from target model outputs. Specifically, we applied spectral graph augmentations to create varied graph perspectives for the surrogate model within a contrastive learning framework at the node level. Thus, the considered method seamlessly aligns with the overarching framework proposed by [30], serving as an enhanced approach for training surrogate models. This adaptation significantly enhances the efficiency and overall performance of the model-stealing process. The conducted comparative analysis against the state of the art showed that the proposed stealing technique demonstrates superior performance while requiring fewer queries of the target model.
@@ Appendix C Additional result - Type I attack
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.764$\pm$0.000 0.852$\pm$0.000 0.772$\pm$0.001 0.889$\pm$0.001 0.768$\pm$0.001 0.855$\pm$0.002
Prediction ours 0.772$\pm$0.001 0.804$\pm$0.001 0.762$\pm$0.003 0.825$\pm$0.004 0.778$\pm$0.002 0.813$\pm$0.002
Shen et al. 0.697$\pm$0.000 0.018$\pm$0.000 0.720$\pm$0.008 0.790$\pm$0.010 0.686$\pm$0.033 0.742$\pm$0.038
Projection ours 0.735$\pm$0.002 0.771$\pm$0.003 0.741$\pm$0.001 0.789$\pm$0.002 0.726$\pm$0.003 0.772$\pm$0.006
Shen et al. 0.759$\pm$0.001 0.846$\pm$0.000 0.776$\pm$0.001 0.872$\pm$0.010 0.761$\pm$0.005 0.863$\pm$0.008
DBLP Embedding ours 0.764$\pm$0.002 0.793$\pm$0.004 0.775$\pm$0.002 0.816$\pm$0.001 0.759$\pm$0.004 0.790$\pm$0.003
Shen et al. 0.840$\pm$0.001 0.912$\pm$0.000 0.824$\pm$0.001 0.931$\pm$0.002 0.838$\pm$0.000 0.917$\pm$0.000
Prediction ours 0.852$\pm$0.001 0.874$\pm$0.001 0.834$\pm$0.001 0.886$\pm$0.002 0.853$\pm$0.001 0.879$\pm$0.001
Shen et al. 0.819$\pm$0.008 0.874$\pm$0.000 0.812$\pm$0.005 0.903$\pm$0.008 0.800$\pm$0.023 0.856$\pm$0.024
Projection ours 0.842$\pm$0.002 0.865$\pm$0.002 0.827$\pm$0.002 0.880$\pm$0.002 0.841$\pm$0.001 0.867$\pm$0.000
Shen et al. 0.843$\pm$0.000 0.915$\pm$0.000 0.827$\pm$0.002 0.919$\pm$0.002 0.838$\pm$0.001 0.930$\pm$0.001
Pubmed Embedding ours 0.854$\pm$0.002 0.871$\pm$0.000 0.836$\pm$0.002 0.880$\pm$0.003 0.854$\pm$0.001 0.859$\pm$0.001
Shen et al. 0.821$\pm$0.002 0.892$\pm$0.000 0.827$\pm$0.003 0.910$\pm$0.001 0.825$\pm$0.003 0.884$\pm$0.002
Prediction ours 0.820$\pm$0.003 0.849$\pm$0.003 0.835$\pm$0.003 0.863$\pm$0.005 0.820$\pm$0.001 0.855$\pm$0.002
Shen et al. 0.626$\pm$0.028 0.645$\pm$0.000 0.744$\pm$0.005 0.780$\pm$0.001 0.649$\pm$0.019 0.672$\pm$0.017
Projection ours 0.731$\pm$0.005 0.749$\pm$0.005 0.753$\pm$0.004 0.777$\pm$0.005 0.686$\pm$0.007 0.689$\pm$0.008
Shen et al. 0.825$\pm$0.003 0.903$\pm$0.000 0.843$\pm$0.002 0.912$\pm$0.001 0.816$\pm$0.003 0.882$\pm$0.003
Citeseer Embedding ours 0.819$\pm$0.003 0.842$\pm$0.003 0.842$\pm$0.006 0.866$\pm$0.005 0.826$\pm$0.004 0.839$\pm$0.006
Shen et al. 0.935$\pm$0.000 0.945$\pm$0.000 0.916$\pm$0.002 0.933$\pm$0.004 0.916$\pm$0.003 0.936$\pm$0.004
Prediction ours 0.925$\pm$0.002 0.929$\pm$0.001 0.900$\pm$0.012 0.917$\pm$0.012 0.914$\pm$0.001 0.921$\pm$0.003
Shen et al. 0.671$\pm$0.020 0.664$\pm$0.000 0.717$\pm$0.026 0.726$\pm$0.025 0.683$\pm$0.023 0.683$\pm$0.028
Projection ours 0.784$\pm$0.011 0.785$\pm$0.011 0.756$\pm$0.014 0.764$\pm$0.014 0.711$\pm$0.018 0.708$\pm$0.022
Shen et al. 0.917$\pm$0.000 0.928$\pm$0.000 0.918$\pm$0.001 0.940$\pm$0.001 0.880$\pm$0.004 0.903$\pm$0.004
Amazon Embedding ours 0.915$\pm$0.003 0.921$\pm$0.001 0.899$\pm$0.005 0.907$\pm$0.008 0.890$\pm$0.006 0.895$\pm$0.003
Shen et al. 0.940$\pm$ 0.000 0.954$\pm$ 0.000 0.926$\pm$ 0.002 0.927$\pm$ 0.002 0.942$\pm$ 0.000 0.962$\pm$ 0.001
Prediction ours 0.946$\pm$0.001 0.954$\pm$0.002 0.935$\pm$0.005 0.950$\pm$0.005 0.939$\pm$0.001 0.951$\pm$0.001
Shen et al. 0.792$\pm$ 0.066 0.801$\pm$ 0.069 0.864$\pm$ 0.005 0.874$\pm$ 0.005 0.727$\pm$ 0.038 0.734$\pm$ 0.941
Projection ours 0.905$\pm$0.008 0.912$\pm$0.010 0.898$\pm$0.008 0.904$\pm$0.008 0.883$\pm$0.003 0.894$\pm$0.004
Shen et al. 0.932$\pm$ 0.002 0.945$\pm$ 0.002 0.938$\pm$ 0.002 0.927$\pm$ 0.002 0.935$\pm$ 0.001 0.953$\pm$ 0.001
Coauthor Embedding ours 0.941$\pm$0.002 0.946$\pm$0.002 0.942$\pm$0.002 0.949$\pm$0.003 0.942$\pm$0.003 0.949$\pm$0.002
Shen et al. 0.898$\pm$0.004 0.926$\pm$0.000 0.889$\pm$0.001 0.939$\pm$0.004 0.893$\pm$0.001 0.934$\pm$0.001
Prediction ours 0.901$\pm$0.004 0.922$\pm$0.006 0.896$\pm$0.001 0.935$\pm$0.005 0.901$\pm$0.002 0.923$\pm$0.004
Shen et al. 0.860$\pm$0.010 0.900$\pm$0.000 0.885$\pm$0.004 0.916$\pm$0.006 0.851$\pm$0.029 0.883$\pm$0.020
Projection ours 0.871$\pm$0.006 0.892$\pm$0.006 0.887$\pm$0.001 0.937$\pm$0.002 0.870$\pm$0.014 0.894$\pm$0.019
Shen et al. 0.898$\pm$0.003 0.931$\pm$0.000 0.893$\pm$0.002 0.937$\pm$0.002 0.831$\pm$0.012 0.855$\pm$0.025
ACM Embedding ours 0.870$\pm$0.011 0.862$\pm$0.009 0.894$\pm$0.005 0.928$\pm$0.003 0.879$\pm$0.010 0.881$\pm$0.010
@@
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.776$\pm$0.001 0.899$\pm$0.000 0.779$\pm$0.001 0.901$\pm$0.003 0.777$\pm$0.003 0.899$\pm$0.004
Prediction ours 0.783$\pm$0.001 0.855$\pm$0.001 0.781$\pm$0.002 0.848$\pm$0.001 0.784$\pm$0.001 0.857$\pm$0.003
Shen et al. 0.714$\pm$0.027 0.811$\pm$0.000 0.744$\pm$0.002 0.839$\pm$0.004 0.739$\pm$0.014 0.836$\pm$0.025
Projection ours 0.750$\pm$0.001 0.827$\pm$0.004 0.754$\pm$0.004 0.829$\pm$0.007 0.749$\pm$0.005 0.821$\pm$0.004
Shen et al. 0.773$\pm$0.001 0.878$\pm$0.000 0.781$\pm$0.002 0.879$\pm$0.001 0.783$\pm$0.000 0.903$\pm$0.002
DBLP Embedding ours 0.773$\pm$0.002 0.840$\pm$0.003 0.779$\pm$0.002 0.842$\pm$0.010 0.763$\pm$0.003 0.815$\pm$0.007
Shen et al. 0.859$\pm$0.001 0.966$\pm$0.000 0.847$\pm$0.000 0.956$\pm$0.001 0.854$\pm$0.000 0.964$\pm$0.002
Prediction ours 0.856$\pm$0.002 0.920$\pm$0.005 0.848$\pm$0.001 0.923$\pm$0.002 0.858$\pm$0.001 0.926$\pm$0.002
Shen et al. 0.821$\pm$0.007 0.890$\pm$0.000 0.841$\pm$0.001 0.945$\pm$0.009 0.844$\pm$0.004 0.936$\pm$0.003
Projection ours 0.849$\pm$0.001 0.915$\pm$0.002 0.844$\pm$0.001 0.922$\pm$0.003 0.843$\pm$0.001 0.911$\pm$0.000
Shen et al. 0.863$\pm$0.001 0.964$\pm$0.000 0.848$\pm$0.001 0.958$\pm$0.000 0.853$\pm$0.002 0.967$\pm$0.003
Pubmed Embedding ours 0.861$\pm$0.001 0.925$\pm$0.000 0.852$\pm$0.001 0.929$\pm$0.001 0.858$\pm$0.002 0.905$\pm$0.002
Shen et al. 0.829$\pm$0.002 0.916$\pm$0.000 0.832$\pm$0.002 0.909$\pm$0.002 0.825$\pm$0.002 0.895$\pm$0.004
Prediction ours 0.813$\pm$0.001 0.861$\pm$0.002 0.833$\pm$0.004 0.873$\pm$0.000 0.824$\pm$0.004 0.864$\pm$0.005
Shen et al. 0.679$\pm$0.030 0.711$\pm$0.000 0.732$\pm$0.015 0.769$\pm$0.014 0.648$\pm$0.043 0.680$\pm$0.039
Projection ours 0.731$\pm$0.003 0.760$\pm$0.002 0.744$\pm$0.005 0.784$\pm$0.004 0.659$\pm$0.024 0.691$\pm$0.021
Shen et al. 0.823$\pm$0.002 0.893$\pm$0.000 0.834$\pm$0.002 0.888$\pm$0.003 0.822$\pm$0.001 0.885$\pm$0.002
Citeseer Embedding ours 0.819$\pm$0.000 0.848$\pm$0.003 0.831$\pm$0.003 0.861$\pm$0.000 0.816$\pm$0.009 0.834$\pm$0.007
Shen et al. 0.926$\pm$0.001 0.932$\pm$0.000 0.897$\pm$0.004 0.910$\pm$0.001 0.888$\pm$0.009 0.913$\pm$0.012
Prediction ours 0.920$\pm$0.001 0.921$\pm$0.001 0.909$\pm$0.004 0.911$\pm$0.001 0.908$\pm$0.003 0.907$\pm$0.005
Shen et al. 0.750$\pm$0.015 0.780$\pm$0.000 0.787$\pm$0.029 0.803$\pm$0.023 0.727$\pm$0.008 0.763$\pm$0.016
Projection ours 0.822$\pm$0.015 0.831$\pm$0.015 0.790$\pm$0.010 0.796$\pm$0.007 0.663$\pm$0.026 0.691$\pm$0.024
Shen et al. 0.930$\pm$0.002 0.925$\pm$0.000 0.915$\pm$0.001 0.929$\pm$0.003 0.918$\pm$0.002 0.931$\pm$0.005
Amazon Embedding ours 0.913$\pm$0.005 0.912$\pm$0.004 0.875$\pm$0.014 0.894$\pm$0.011 0.907$\pm$0.005 0.904$\pm$0.005
Shen et al. 0.945$\pm$0.001 0.971$\pm$0.000 0.904$\pm$0.006 0.929$\pm$0.006 0.943$\pm$0.001 0.974$\pm$0.002
Prediction ours 0.942$\pm$0.004 0.962$\pm$0.005 0.934$\pm$0.003 0.956$\pm$0.005 0.941$\pm$0.000 0.968$\pm$0.001
Shen et al. 0.833$\pm$0.029 0.849$\pm$0.000 0.822$\pm$0.028 0.845$\pm$0.030 0.755$\pm$0.021 0.776$\pm$0.017
Projection ours 0.927$\pm$0.000 0.949$\pm$0.001 0.908$\pm$0.006 0.930$\pm$0.006 0.905$\pm$0.006 0.927$\pm$0.007
Shen et al. 0.940$\pm$0.001 0.964$\pm$0.000 0.886$\pm$0.008 0.902$\pm$0.009 0.939$\pm$0.001 0.969$\pm$0.002
Coauthor Embedding ours 0.942$\pm$0.002 0.959$\pm$0.003 0.942$\pm$0.002 0.954$\pm$0.004 0.940$\pm$0.003 0.950$\pm$0.003
Shen et al. 0.898$\pm$0.003 0.933$\pm$0.000 0.891$\pm$0.003 0.941$\pm$0.003 0.895$\pm$0.003 0.935$\pm$0.003
Prediction ours 0.888$\pm$0.003 0.903$\pm$0.006 0.892$\pm$0.003 0.923$\pm$0.017 0.888$\pm$0.004 0.914$\pm$0.004
Shen et al. 0.849$\pm$0.047 0.884$\pm$0.100 0.869$\pm$0.002 0.913$\pm$0.002 0.891$\pm$0.005 0.931$\pm$0.008
Projection ours 0.887$\pm$0.004 0.912$\pm$0.003 0.887$\pm$0.008 0.919$\pm$0.010 0.865$\pm$0.014 0.887$\pm$0.017
Shen et al. 0.902$\pm$0.002 0.936$\pm$0.000 0.876$\pm$0.001 0.925$\pm$0.002 0.861$\pm$0.014 0.889$\pm$0.009
ACM Embedding ours 0.883$\pm$0.007 0.891$\pm$0.010 0.882$\pm$0.008 0.911$\pm$0.015 0.870$\pm$0.007 0.874$\pm$0.013

### v4

@@ 8 Conclusions
In this work, we identify a new unsupervised GNN model-stealing attack. Only a single work [33] targeting inductive GNNs existed previously. Our proposed approach combines a contrastive learning objective with graph-specific augmentations, improving the efficient utilization of information derived from target model outputs. Specifically, we applied spectral graph augmentations to create varied graph perspectives for the surrogate model within a contrastive learning framework at the node level. Thus, the considered method seamlessly aligns with the overarching framework proposed by [33], serving as an enhanced approach for training surrogate models. This adaptation significantly enhances the efficiency and overall performance of the model-stealing process. The conducted comparative analysis against the state of the art showed that the proposed stealing technique demonstrates superior performance while requiring fewer queries of the target model.
@@ Appendix C Additional result - Type I attack
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.764$\pm$0.000 0.852$\pm$0.000 0.695$\pm$0.001 0.772$\pm$0.001 0.889$\pm$0.001 0.688$\pm$0.005 0.768$\pm$0.001 0.855$\pm$0.002 0.698$\pm$0.005
Prediction ours 0.772$\pm$0.001 0.804$\pm$0.001 0.706$\pm$0.009 0.762$\pm$0.003 0.825$\pm$0.004 0.692$\pm$0.001 0.778$\pm$0.002 0.813$\pm$0.002 0.691$\pm$0.003
Shen et al. 0.697$\pm$0.000 0.618$\pm$0.000 0.527$\pm$0.004 0.720$\pm$0.008 0.790$\pm$0.010 0.604$\pm$0.003 0.686$\pm$0.033 0.742$\pm$0.038 0.363$\pm$0.041
Projection ours 0.735$\pm$0.002 0.771$\pm$0.003 0.609$\pm$0.001 0.741$\pm$0.001 0.789$\pm$0.002 0.625$\pm$0.001 0.726$\pm$0.003 0.772$\pm$0.006 0.541$\pm$0.005
Shen et al. 0.759$\pm$0.001 0.846$\pm$0.000 0.673$\pm$0.009 0.776$\pm$0.001 0.872$\pm$0.010 0.706$\pm$0.001 0.761$\pm$0.005 0.863$\pm$0.008 0.651$\pm$0.005
DBLP Embedding ours 0.764$\pm$0.002 0.793$\pm$0.004 0.683$\pm$0.009 0.775$\pm$0.002 0.816$\pm$0.001 0.699$\pm$0.002 0.759$\pm$0.004 0.790$\pm$0.003 0.696$\pm$0.001
Shen et al. 0.840$\pm$0.001 0.912$\pm$0.000 0.836$\pm$0.002 0.824$\pm$0.001 0.931$\pm$0.002 0.819$\pm$0.003 0.838$\pm$0.000 0.917$\pm$0.000 0.832$\pm$0.003
Prediction ours 0.852$\pm$0.001 0.874$\pm$0.001 0.847$\pm$0.001 0.834$\pm$0.001 0.886$\pm$0.002 0.836$\pm$0.004 0.853$\pm$0.001 0.879$\pm$0.001 0.848$\pm$0.003
Shen et al. 0.819$\pm$0.008 0.874$\pm$0.000 0.722$\pm$0.004 0.812$\pm$0.005 0.903$\pm$0.008 0.799$\pm$0.005 0.800$\pm$0.023 0.856$\pm$0.024 0.809$\pm$0.001
Projection ours 0.842$\pm$0.002 0.865$\pm$0.002 0.833$\pm$0.001 0.827$\pm$0.002 0.880$\pm$0.002 0.815$\pm$0.001 0.841$\pm$0.001 0.867$\pm$0.000 0.827$\pm$0.003
Shen et al. 0.843$\pm$0.000 0.915$\pm$0.000 0.839$\pm$0.001 0.827$\pm$0.002 0.919$\pm$0.002 0.822$\pm$0.005 0.838$\pm$0.001 0.930$\pm$0.001 0.833$\pm$0.009
Pubmed Embedding ours 0.854$\pm$0.002 0.871$\pm$0.000 0.852$\pm$0.009 0.836$\pm$0.002 0.880$\pm$0.003 0.831$\pm$0.003 0.854$\pm$0.001 0.859$\pm$0.001 0.857$\pm$0.004
Shen et al. 0.821$\pm$0.002 0.892$\pm$0.000 0.824$\pm$0.005 0.827$\pm$0.003 0.910$\pm$0.001 0.827$\pm$0.001 0.825$\pm$0.003 0.884$\pm$0.002 0.830$\pm$0.003
Prediction ours 0.820$\pm$0.003 0.849$\pm$0.003 0.821$\pm$0.002 0.835$\pm$0.003 0.863$\pm$0.005 0.842$\pm$0.001 0.820$\pm$0.001 0.855$\pm$0.002 0.822$\pm$0.005
Shen et al. 0.626$\pm$0.028 0.645$\pm$0.000 0.657$\pm$0.001 0.744$\pm$0.005 0.780$\pm$0.001 0.753$\pm$0.001 0.649$\pm$0.019 0.672$\pm$0.017 0.694$\pm$0.001
Projection ours 0.731$\pm$0.005 0.749$\pm$0.005 0.724$\pm$0.001 0.753$\pm$0.004 0.777$\pm$0.005 0.747$\pm$0.001 0.686$\pm$0.007 0.689$\pm$0.008 0.696$\pm$0.004
Shen et al. 0.825$\pm$0.003 0.903$\pm$0.000 0.821$\pm$0.009 0.843$\pm$0.002 0.912$\pm$0.001 0.845$\pm$0.002 0.816$\pm$0.003 0.882$\pm$0.003 0.813$\pm$0.004
Citeseer Embedding ours 0.819$\pm$0.003 0.842$\pm$0.003 0.823$\pm$0.005 0.842$\pm$0.006 0.866$\pm$0.005 0.848$\pm$0.005 0.826$\pm$0.004 0.839$\pm$0.006 0.824$\pm$0.001
Shen et al. 0.935$\pm$0.000 0.945$\pm$0.000 0.923$\pm$0.001 0.916$\pm$0.002 0.933$\pm$0.004 0.906$\pm$0.009 0.916$\pm$0.003 0.936$\pm$0.004 0.897$\pm$0.002
Prediction ours 0.925$\pm$0.002 0.929$\pm$0.001 0.912$\pm$0.003 0.900$\pm$0.012 0.917$\pm$0.012 0.853$\pm$0.009 0.914$\pm$0.001 0.921$\pm$0.003 0.877$\pm$0.009
Shen et al. 0.671$\pm$0.020 0.664$\pm$0.000 0.552$\pm$0.002 0.717$\pm$0.026 0.726$\pm$0.025 0.592$\pm$0.012 0.683$\pm$0.023 0.683$\pm$0.028 0.549$\pm$0.003
Projection ours 0.784$\pm$0.011 0.785$\pm$0.011 0.708$\pm$0.005 0.756$\pm$0.014 0.764$\pm$0.014 0.681$\pm$0.001 0.711$\pm$0.018 0.708$\pm$0.022 0.572$\pm$0.003
Shen et al. 0.917$\pm$0.000 0.928$\pm$0.000 0.892$\pm$0.009 0.918$\pm$0.001 0.940$\pm$0.001 0.908$\pm$0.002 0.880$\pm$0.004 0.903$\pm$0.004 0.850$\pm$0.003
Amazon Embedding ours 0.915$\pm$0.003 0.921$\pm$0.001 0.889$\pm$0.004 0.899$\pm$0.005 0.907$\pm$0.008 0.883$\pm$0.003 0.890$\pm$0.006 0.895$\pm$0.003 0.851$\pm$0.002
Shen et al. 0.940$\pm$ 0.000 0.954$\pm$0.000 0.919$\pm$0.005 0.926$\pm$ 0.002 0.927$\pm$ 0.002 0.908$\pm$0.001 0.942$\pm$ 0.000 0.962$\pm$ 0.001 0.922$\pm$0.009
Prediction ours 0.946$\pm$0.001 0.954$\pm$0.002 0.929$\pm$0.004 0.935$\pm$0.005 0.950$\pm$0.005 0.905$\pm$0.004 0.939$\pm$0.001 0.951$\pm$0.001 0.919$\pm$0.002
Shen et al. 0.792$\pm$ 0.066 0.801$\pm$ 0.069 0.593$\pm$0.001 0.864$\pm$ 0.005 0.874$\pm$ 0.005 0.823$\pm$0.001 0.727$\pm$ 0.038 0.734$\pm$ 0.941 0.650$\pm$0.003
Projection ours 0.905$\pm$0.008 0.912$\pm$0.010 0.847$\pm$0.001 0.898$\pm$0.008 0.904$\pm$0.008 0.868$\pm$0.001 0.883$\pm$0.003 0.894$\pm$0.004 0.827$\pm$0.004
Shen et al. 0.932$\pm$ 0.002 0.945$\pm$ 0.002 0.910$\pm$0.005 0.938$\pm$ 0.002 0.927$\pm$ 0.002 0.916$\pm$0.001 0.935$\pm$ 0.001 0.953$\pm$ 0.001 0.912$\pm$0.002
Coauthor Embedding ours 0.941$\pm$0.002 0.946$\pm$0.002 0.922$\pm$0.001 0.942$\pm$0.002 0.949$\pm$0.003 0.927$\pm$0.009 0.942$\pm$0.003 0.949$\pm$0.002 0.923$\pm$0.003
Shen et al. 0.898$\pm$0.004 0.926$\pm$0.000 0.896$\pm$0.004 0.889$\pm$0.001 0.939$\pm$0.004 0.883$\pm$0.002 0.893$\pm$0.001 0.934$\pm$0.001 0.895$\pm$0.001
Prediction ours 0.901$\pm$0.004 0.922$\pm$0.006 0.898$\pm$0.002 0.896$\pm$0.001 0.935$\pm$0.005 0.895$\pm$0.002 0.901$\pm$0.002 0.923$\pm$0.004 0.900$\pm$0.009
Shen et al. 0.860$\pm$0.010 0.900$\pm$0.000 0.765$\pm$0.002 0.885$\pm$0.004 0.916$\pm$0.006 0.881$\pm$0.002 0.851$\pm$0.029 0.883$\pm$0.020 0.873$\pm$0.001
Projection ours 0.871$\pm$0.006 0.892$\pm$0.006 0.883$\pm$0.003 0.887$\pm$0.001 0.937$\pm$0.002 0.882$\pm$0.009 0.870$\pm$0.014 0.894$\pm$0.019 0.898$\pm$0.005
Shen et al. 0.898$\pm$0.003 0.931$\pm$0.000 0.887$\pm$0.002 0.893$\pm$0.002 0.937$\pm$0.002 0.892$\pm$0.005 0.831$\pm$0.012 0.855$\pm$0.025 0.833$\pm$0.004
ACM Embedding ours 0.870$\pm$0.011 0.862$\pm$0.009 0.877$\pm$0.001 0.894$\pm$0.005 0.928$\pm$0.003 0.899$\pm$0.004 0.879$\pm$0.010 0.881$\pm$0.010 0.871$\pm$0.004
@@
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.776$\pm$0.001 0.899$\pm$0.000 0.698$\pm$0.001 0.779$\pm$0.001 0.901$\pm$0.003 0.694$\pm$0.004 0.777$\pm$0.003 0.899$\pm$0.004 0.697$\pm$0.001
Prediction ours 0.783$\pm$0.001 0.855$\pm$0.001 0.712$\pm$0.003 0.781$\pm$0.002 0.848$\pm$0.001 0.734$\pm$0.003 0.784$\pm$0.001 0.857$\pm$0.003 0.683$\pm$0.001
Shen et al. 0.714$\pm$0.027 0.811$\pm$0.000 0.557$\pm$0.002 0.744$\pm$0.002 0.839$\pm$0.004 0.604$\pm$0.005 0.739$\pm$0.014 0.836$\pm$0.025 0.484$\pm$0.012
Projection ours 0.750$\pm$0.001 0.827$\pm$0.004 0.618$\pm$0.001 0.754$\pm$0.004 0.829$\pm$0.007 0.620$\pm$0.003 0.749$\pm$0.005 0.821$\pm$0.004 0.616$\pm$0.005
Shen et al. 0.773$\pm$0.001 0.878$\pm$0.000 0.700$\pm$0.005 0.781$\pm$0.002 0.879$\pm$0.001 0.715$\pm$0.001 0.783$\pm$0.000 0.903$\pm$0.002 0.706$\pm$0.003
DBLP Embedding ours 0.773$\pm$0.002 0.840$\pm$0.003 0.715$\pm$0.009 0.779$\pm$0.002 0.842$\pm$0.010 0.706$\pm$0.001 0.763$\pm$0.003 0.815$\pm$0.007 0.707$\pm$0.001
Shen et al. 0.859$\pm$0.001 0.966$\pm$0.000 0.856$\pm$0.002 0.847$\pm$0.000 0.956$\pm$0.001 0.845$\pm$0.005 0.854$\pm$0.000 0.964$\pm$0.002 0.852$\pm$0.001
Pre diction ours 0.856$\pm$0.002 0.920$\pm$0.005 0.851$\pm$0.003 0.848$\pm$0.001 0.923$\pm$0.002 0.849$\pm$0.005 0.858$\pm$0.001 0.926$\pm$0.002 0.854$\pm$0.005
Shen et al. 0.821$\pm$0.007 0.890$\pm$0.000 0.757$\pm$0.005 0.841$\pm$0.001 0.945$\pm$0.009 0.823$\pm$0.004 0.844$\pm$0.004 0.936$\pm$0.003 0.832$\pm$0.003
Projection ours 0.849$\pm$0.001 0.915$\pm$0.002 0.842$\pm$0.003 0.844$\pm$0.001 0.922$\pm$0.003 0.836$\pm$0.004 0.843$\pm$0.001 0.911$\pm$0.000 0.825$\pm$0.003
Shen et al. 0.863$\pm$0.001 0.964$\pm$0.000 0.860$\pm$0.001 0.848$\pm$0.001 0.958$\pm$0.000 0.847$\pm$0.002 0.853$\pm$0.002 0.967$\pm$0.003 0.852$\pm$0.009
Pubmed Embedding ours 0.861$\pm$0.001 0.925$\pm$0.000 0.854$\pm$0.001 0.852$\pm$0.001 0.929$\pm$0.001 0.849$\pm$0.005 0.858$\pm$0.002 0.905$\pm$0.002 0.846$\pm$0.009
Shen et al. 0.829$\pm$0.002 0.916$\pm$0.000 0.831$\pm$0.003 0.832$\pm$0.002 0.909$\pm$0.002 0.829$\pm$0.005 0.825$\pm$0.002 0.895$\pm$0.004 0.902$\pm$0.005
Prediction ours 0.813$\pm$0.001 0.861$\pm$0.002 0.809$\pm$0.009 0.833$\pm$0.004 0.873$\pm$0.000 0.833$\pm$0.009 0.824$\pm$0.004 0.864$\pm$0.005 0.816$\pm$0.009
Shen et al. 0.679$\pm$0.030 0.711$\pm$0.000 0.554$\pm$0.001 0.732$\pm$0.015 0.769$\pm$0.014 0.692$\pm$0.005 0.648$\pm$0.043 0.680$\pm$0.039 0.658$\pm$0.022
Projection ours 0.731$\pm$0.003 0.760$\pm$0.002 0.723$\pm$0.004 0.744$\pm$0.005 0.784$\pm$0.004 0.738$\pm$0.005 0.659$\pm$0.024 0.691$\pm$0.021 0.686$\pm$0.023
Shen et al. 0.823$\pm$0.002 0.893$\pm$0.000 0.823$\pm$0.001 0.834$\pm$0.002 0.888$\pm$0.003 0.831$\pm$0.005 0.822$\pm$0.001 0.885$\pm$0.002 0.892$\pm$0.001
Citeseer Embedding ours 0.819$\pm$0.000 0.848$\pm$0.003 0.819$\pm$0.009 0.831$\pm$0.003 0.861$\pm$0.000 0.827$\pm$0.003 0.816$\pm$0.009 0.834$\pm$0.007 0.813$\pm$0.005
Shen et al. 0.926$\pm$0.001 0.932$\pm$0.000 0.910$\pm$0.009 0.897$\pm$0.004 0.910$\pm$0.001 0.882$\pm$0.003 0.888$\pm$0.009 0.913$\pm$0.012 0.876$\pm$0.009
Prediction ours 0.920$\pm$0.001 0.921$\pm$0.001 0.899$\pm$0.001 0.909$\pm$0.004 0.911$\pm$0.001 0.879$\pm$0.009 0.908$\pm$0.003 0.907$\pm$0.005 0.863$\pm$0.003
Shen et al. 0.750$\pm$0.015 0.780$\pm$0.000 0.597$\pm$0.001 0.787$\pm$0.029 0.803$\pm$0.023 0.704$\pm$0.005 0.727$\pm$0.008 0.763$\pm$0.016 0.567$\pm$0.005
Projection ours 0.822$\pm$0.015 0.831$\pm$0.015 0.671$\pm$0.002 0.790$\pm$0.010 0.796$\pm$0.007 0.648$\pm$0.009 0.663$\pm$0.026 0.691$\pm$0.024 0.543$\pm$0.009
Shen et al. 0.930$\pm$0.002 0.925$\pm$0.000 0.920$\pm$0.005 0.915$\pm$0.001 0.929$\pm$0.003 0.893$\pm$0.001 0.918$\pm$0.002 0.931$\pm$0.005 0.904$\pm$0.002
Amazon Embedding ours 0.913$\pm$0.005 0.912$\pm$0.004 0.887$\pm$0.003 0.875$\pm$0.014 0.894$\pm$0.011 0.844$\pm$0.001 0.907$\pm$0.005 0.904$\pm$0.005 0.876$\pm$0.001
Shen et al. 0.945$\pm$0.001 0.971$\pm$0.000 0.924$\pm$0.002 0.904$\pm$0.006 0.929$\pm$0.006 0.891$\pm$0.009 0.943$\pm$0.001 0.974$\pm$0.002 0.925$\pm$0.004
Prediction ours 0.942$\pm$0.004 0.962$\pm$0.005 0.922$\pm$0.001 0.934$\pm$0.003 0.956$\pm$0.005 0.901$\pm$0.004 0.941$\pm$0.000 0.968$\pm$0.001 0.926$\pm$0.009
Shen et al. 0.833$\pm$0.029 0.849$\pm$0.000 0.553$\pm$0.004 0.822$\pm$0.028 0.845$\pm$0.030 0.684$\pm$0.001 0.755$\pm$0.021 0.776$\pm$0.017 0.456$\pm$0.031
Projection ours 0.927$\pm$0.000 0.949$\pm$0.001 0.900$\pm$0.001 0.908$\pm$0.006 0.930$\pm$0.006 0.860$\pm$0.005 0.905$\pm$0.006 0.927$\pm$0.007 0.863$\pm$0.002
Shen et al. 0.940$\pm$0.001 0.964$\pm$0.000 0.918$\pm$0.001 0.886$\pm$0.008 0.902$\pm$0.009 0.901$\pm$0.009 0.939$\pm$0.001 0.969$\pm$0.002 0.921$\pm$0.002
Coauthor Embedding ours 0.942$\pm$0.002 0.959$\pm$0.003 0.926$\pm$0.001 0.942$\pm$0.002 0.954$\pm$0.004 0.918$\pm$0.005 0.940$\pm$0.003 0.950$\pm$0.003 0.923$\pm$0.004
Shen et al. 0.898$\pm$0.003 0.933$\pm$0.000 0.899$\pm$0.002 0.891$\pm$0.003 0.941$\pm$0.003 0.889$\pm$0.003 0.895$\pm$0.003 0.935$\pm$0.003 0.883$\pm$0.005
Prediction ours 0.888$\pm$0.003 0.903$\pm$0.006 0.887$\pm$0.005 0.892$\pm$0.003 0.923$\pm$0.017 0.896$\pm$0.009 0.888$\pm$0.004 0.914$\pm$0.004 0.892$\pm$0.001
Shen et al. 0.849$\pm$0.047 0.884$\pm$0.100 0.887$\pm$0.003 0.869$\pm$0.002 0.913$\pm$0.002 0.863$\pm$0.002 0.891$\pm$0.005 0.931$\pm$0.008 0.888$\pm$0.001
Projection ours 0.887$\pm$0.004 0.912$\pm$0.003 0.892$\pm$0.004 0.887$\pm$0.008 0.919$\pm$0.010 0.892$\pm$0.004 0.865$\pm$0.014 0.887$\pm$0.017 0.887$\pm$0.001
Shen et al. 0.902$\pm$0.002 0.936$\pm$0.000 0.886$\pm$0.004 0.876$\pm$0.001 0.925$\pm$0.002 0.866$\pm$0.001 0.861$\pm$0.014 0.889$\pm$0.009 0.852$\pm$0.002
ACM Embedding ours 0.883$\pm$0.007 0.891$\pm$0.010 0.873$\pm$0.001 0.882$\pm$0.008 0.911$\pm$0.015 0.886$\pm$0.001 0.870$\pm$0.007 0.874$\pm$0.013 0.870$\pm$0.004